# MOGO: Residual Quantized Hierarchical Causal Transformer for High-Quality and Real-Time 3D Human Motion Generation

### v1

@@ 3.3.2 Quantitative Results
| MOGO† | 0.5150.003 | 0.7090.003 | 0.8010.003 | 0.0380.002 | 2.9510.008 | 2.1080.070 | |
| MOGO with TCA† | 0.5270.007 | 0.7220.008 | 0.8270.012 | 0.0640.003 | 2.8490.003 | 2.3440.037 | |
@@
| MOGO† | 0.4200.007 | 0.6340.007 | 0.7540.007 | 0.1910.016 | 2.9570.029 | 2.0630.066 | |
| MOGO with TCA† | 0.4470.023 | 0.6680.016 | 0.8010.007 | 0.3130.009 | 2.8490.007 | 2.2730.073 | |
@@
Red = Best, Blue = Second Best, orange = Third Best.
This advantage becomes more pronounced in the CMP zero-shot setting, where other transformer-based approaches often suffer from significant degradation. In contrast, MOGO with TCA achieves the best results in multiple metrics, highlighting its ability to generalize effectively to out-of-distribution prompts. These improvements stem from the synergy between our hierarchical generation framework and the inference-time input alignment mechanism introduced by TCA.
Furthermore, MOGO supports streaming token-wise generation, allowing for real-time inference without sacrificing output quality. This combination of strong generalization, high overall performance, and low-latency generation makes MOGO well-suited for practical deployment in open-ended, real-world motion generation scenarios.
3.3.3 Qualitative Comparison and Long-range Generation
Qualitative Comparison. As shown in Figure 3, our generated motions are smoother, more consistent and better aligned with input prompts compared to other methods (more results are shown in Appendix A, and videos are also provided in the supplementary material). Built upon a hierarchical architecture, MOGO enables longer motion generation without sacrificing coherence or realism, demonstrating competitive performance among transformer-based autoregressive models and strong generalization capabilities (more ablation studies are shown in Appendix).
4 Related Work
LLM-Based Motion Generation Models. LLM-based architectures have become a cornerstone of text-to-motion generation, leveraging their ability to model sequential data and adapt to varied tasks [2, 3, 4, 5, 20]. Models like MoMask [2] and MMM [4] use masked token modeling to produce high-quality motions but struggle with streaming output and generalization in low-data or out-of-distribution settings due to their bidirectional design. Conversely, autoregressive models like T2M-GPT [3] and AttT2M [5] enable sequential generation, making them suitable for real-time applications and scalable with larger datasets, though they often sacrifice some motion quality. Efforts like MotionGPT [20] integrate multimodal language modeling but face challenges in achieving high-fidelity motion outputs. Our MOGO framework addresses these issues by combining efficient encoding through the MoSA-VQ with a single-pass autoregressive transformer, ensuring both high-quality motion and streaming capabilities.
5 Limitations and Discussion
Despite MOGO’s strong performance in real-time streaming and zero-shot generalization, several limitations remain.
Motion Editing. Due to its unidirectional autoregressive design, MOGO struggles with temporal editing tasks such as infilling or interpolation. This limits its use in scenarios like animation post-processing, where flexible motion manipulation is needed.
Generation Length. While relative positional encoding allows MOGO to generate up to 260 frames in cyclic motions, non-continuous sequences remain limited to around 196 frames. This is primarily due to weak temporal correlations in motion data, which hinder the use of segment-level memory mechanisms like Transformer-XL [34].
To address these issues, future work should focus on (1) enabling conditional adjustments after initial generation to support dynamic prompts, and (2) improving local condition alignment for seamless mid-sequence control. These enhancements could make MOGO more adaptable for interactive and editable motion generation.
6 Conclusion
@@
We presented MOGO, a one-pass autoregressive framework for high-quality and real-time 3D human motion generation. By combining the MoSA-VQ module for scale-adaptive residual vector quantization with the RQHC-Transformer for hierarchical causal decoding, MOGO achieves efficient and temporally coherent motion synthesis. The integration of TCA mechanism further enhances semantic alignment and enables better generalization in zero- and few-shot scenarios. Our framework supports low-latency, frame-by-frame inference, making it well-suited for real-time and streamable applications. Extensive experiments demonstrate that MOGO outperforms prior transformer-based methods across multiple benchmarks in terms of generation quality and text-motion alignment. While current limitations remain in motion editing and handling complex non-cyclic sequences, MOGO provides a robust foundation for future research toward controllable, editable, and interactive motion generation.
Appendix A Generation Results
Stylized Motion Generation.
Figure 4 illustrates the results of our model applied to stylized motion generation. Three representative action types are shown, each demonstrating the model’s ability to preserve motion semantics while adapting to distinct style patterns. To further showcase the generality and diversity of our approach, we provide additional stylized motion sequences in MP4 format in the supplementary material.
Qualitative Comparison.
Figures 6–9 present a qualitative comparison of motion generation results given the same text prompt, showcasing outputs among our model, Momask [2] and MotionGPT [20]. As observed, MOGO produces more complete and coherent motion sequences, particularly excelling in motion completeness and fidelity. For a more intuitive evaluation, we additionally provide corresponding video results in the supplementary material (MP4 format), which further highlight the differences in motion quality and realism.
Appendix B Ablation Study
Impact of Learnable Scaling in MoSA-VQ.
We investigate the role of the proposed learnable scaling mechanism in MoSA-VQ by comparing models with and without this design. In the baseline configuration, residual features at each quantization level are directly quantized without any adaptive normalization. In contrast, our full model applies a layer-specific affine transformation before quantization, allowing each level to adjust the residual space dynamically.
@@ Appendix B Ablation Study
As shown in Table 3, incorporating learnable scale and bias parameters leads to consistent improvements across multiple evaluation metrics. The model demonstrates better reconstruction quality, enhanced alignment between motion and text, and stronger semantic consistency. These results highlight the benefit of introducing adaptive normalization in residual quantization, which facilitates more expressive and stable token representations.
| Codebook size | w/ Learnable Scaling | ||
| FID | Top1 | MM-Dist | |
@@
| Codebook size | w/o Learnable Scaling | ||
| FID | Top1 | MM-Dist | |
@@
Effect of Quantization Depth in MoSA-VQ.
We further explore how the number of residual quantization layers in MoSA-VQ affects the model’s overall performance. As residual quantization hierarchically captures motion information from coarse to fine granularity, increasing the depth allows more detailed refinement of the motion representation.
| Layers | Reconstruction | Generation | ||||
| FID | Top1 | MM-Dist | FID | Top1 | MM-Dist | |
| 1 | 0.0670.001 | 0.5080.002 | 2.9970.007 | 0.0680.009 | 0.4690.002 | 3.1380.006 |
| 2 | 0.0460.001 | 0.5040.001 | 2.9840.010 | 0.0530.008 | 0.4670.003 | 3.1140.002 |
| 3 | 0.0350.001 | 0.5080.003 | 2.9800.004 | 0.0440.005 | 0.4980.002 | 2.9970.008 |
| 4 | 0.0230.001 | 0.5050.003 | 2.9890.006 | 0.0450.005 | 0.5010.002 | 2.9820.002 |
| 5 | 0.0170.001 | 0.5110.001 | 2.9800.003 | 0.0410.005 | 0.4980.002 | 2.9770.002 |
| 6 | 0.0130.001 | 0.5100.001 | 2.9820.006 | 0.0380.003 | 0.5150.002 | 2.9510.003 |
| 7 | 0.0130.001 | 0.5030.002 | 2.9930.007 | 0.0400.003 | 0.5130.007 | 2.9430.003 |
As shown in Table 4, models with a shallow quantization depth tend to underperform, suggesting that limited hierarchy constrains the representational capacity of the latent code. As the number of layers increases, performance improves consistently across reconstruction fidelity, text-motion alignment, and multimodal distance. However, the gains tend to plateau beyond a certain depth, indicating diminishing returns. This analysis supports the use of a moderate number of residual layers as a trade-off between expressiveness and efficiency.
Codebook Size. We evaluate various codebook sizes on HumanML3D. As shown in Table 5, achieves the best FID for generation and competitive reconstruction.
Dataset Size. We evaluate MOGO’s scalability by mixing HumanML3D with varying proportions of CMP data (Table 6). Without retraining the MoSA-VQ, generation quality improves significantly with increased data volume.

### v2

@@ III-E Text Condition Alignment for Improved Motion Decoding
| MotionAnything [30] | 0.5460.002 | 0.7350.002 | 0.8290.002 | 0.0280.001 | 2.8590.010 | 2.7050.060 | |
| MOGO† | 0.5150.003 | 0.7090.003 | 0.8010.003 | 0.0640.002 | 2.9510.008 | 2.1080.070 | |
| MOGO with TCA† | 0.5270.007 | 0.7220.008 | 0.8270.012 | 0.0380.003 | 2.8490.003 | 2.3440.037 | |
@@
| MotionAnything [30] | 0.4490.007 | 0.6780.004 | 0.8020.006 | 0.1310.003 | 2.7050.024 | 1.3740.060 | |
| MOGO† | 0.4200.007 | 0.6340.007 | 0.7540.007 | 0.3130.016 | 2.9570.029 | 2.0630.066 | |
| MOGO with TCA† | 0.4470.023 | 0.6680.016 | 0.8010.007 | 0.1910.009 | 2.8490.007 | 2.2730.073 | |
@@
IV Experiments
IV-A Datasets and Evaluation Metrics
IV-A1 Datasets
We train and evaluate MOGO on two widely used benchmarks, HumanML3D [17] and KIT-ML [31], following the data splits of T2M [17]. To assess out-of-distribution generalization, we further evaluate zero-shot performance on CMP [32], which contains non-daily, combat-style motions. Prompt engineering is disabled during CMP evaluation to ensure fairness.
IV-A2 Evaluation Metrics
We adopt standard metrics from [17]. FID (Frechet Inception Distance) measures motion realism and serves as a primary evaluation metric. R-Precision evaluates text-motion alignment (Top-1, 2, 3). MM-Dist computes the distance between motion and text embeddings. MultiModality measures the variance of motions generated from the same text prompt.
IV-A3 Training Details
The codebook is sized at with 6 quantization layers, using stride-1 convolutions and 0.2 dropout. It is trained with AdamW (learning rate , batch size 512) for 2000 iterations on an NVIDIA 4090 GPU. RQHC-Transformer has 6 sub-modules aligned with RVQ layers, with head counts and layer counts , and a model dimension of 1024. For HumanML3D, it is trained on an A800-80G with cosine decay from to , batch size 32, for 1500 epochs. For KIT-ML, a V100-32G is used with a learning rate from to , batch size 48.
IV-B Comparison to State-of-the-art Approaches
As shown in Figure 2, motions produced by our model are more semantically consistent with the input descriptions and exhibit smoother, more stable transitions without noticeable spatial drift. Additional qualitative visualizations, including extended motion generation results, are provided in the Appendix.
| HumanML3D | KIT-ML | ||
| Method | FID | Method | FID |
| M2DM | 0.063±0.001 | M2DM | 0.413±0.009 |
| T2M-GPT | 0.070±0.001 | T2M-GPT | 0.472±0.011 |
| MoMask | 0.019±0.001 | MoMask | 0.112±0.002 |
| MMM | 0.075±0.001 | MMM | 0.641±0.014 |
| MOGO (ours) | 0.013±0.001 | MOGO (ours) | 0.037±0.001 |
IV-B1 Reconstruction Quality of Motion Encoder
We compare our VAE design with existing motion VAEs—M2DM, T2M-GPT, MoMask, and MMM—on HumanML3D and KIT-ML (Table II). MOGO achieves the best results on both benchmarks, with an FID of 0.013 on HumanML3D, outperforming MoMask (0.019) and others. On the more challenging KIT-ML dataset, MOGO achieves an FID of 0.037, significantly lower than MoMask (0.112), T2M-GPT (0.472), and MMM (0.641). These results demonstrate the strong reconstruction capability of our VAE across diverse motion distributions.
IV-B2 Quantitative Results.
MOGO demonstrates strong performance across HumanML3D, KIT-ML, and zero-shot CMP benchmarks, particularly on FID and R@3. As showned in Table I, On HumanML3D, our base model achieves an FID of 0.038, outperforming T2M-GPT (0.116), MotionGPT (0.232), and MotionDiffuse (0.630). With TCA, FID slightly increases to 0.064, while R@3 improves from 0.801 to 0.827. On KIT-ML, MOGO attains an FID of 0.191 and R@3 of 0.801, competitive with MotionAnything (FID 0.131, R@3 0.802). Under the zero-shot CMP setting, FID drops to 10.388 (6.873 with TCA), outperforming MotionGPT (15.654) and T2M-GPT (16.092), while R@3 reaches 0.304, over twice that of the best baseline MMM (0.154).
Overall, our method consistently achieves top FID and R@3 scores across benchmarks. The TCA module further enhances semantic alignment, especially under distribution shifts, validating the effectiveness of structured condition interpretation.
IV-C Ablation Study
IV-C1 Ablation Study on VAE Depth
| Depth | FID | R TOP1 | MM-Dist |
| 1 | 0.070±0.001 | 0.502 ±0.001 | 2.999 ±0.006 |
| 3 | 0.021±0.001 | 0.508 ±0.001 | 2.992 ±0.008 |
| 6 | 0.016±0.001 | 0.510 ±0.001 | 2.989 ±0.007 |
| 7 | 0.016±0.001 | 0.509 ±0.001 | 2.996 ±0.003 |
We evaluate the impact of quantization depth by varying the number of layers while fixing the codebook size to (Table VIII). A single-layer quantizer performs poorly (FID: 0.070), indicating limited capacity. Increasing to 3 layers significantly improves performance (FID: 0.021), and 6 layers yields the best overall results (FID: 0.016, R@1: 0.510, MM-Dist: 2.989). Adding a 7th layer offers no further gain and slightly worsens MM-Dist, suggesting diminishing returns. Overall, 6-layer quantization provides the best balance between expressiveness and generalization.
IV-C2 Ablation Study on Layer-Head Configuration
| Heads | Layers | FID | R@1 | MM-Dist |
| [4, 4, 4, 4, 4, 4] | [6] | 0.242±0.009 | 0.397±0.005 | 4.132±0.013 |
| [4, 4, 4, 4, 4, 4] | [6, 6, 6, 6, 6, 6] | 0.101±0.004 | 0.461±0.003 | 3.359±0.009 |
| [12, 6, 4, 2, 2, 2] | [16, 8, 6, 4, 2, 2] | 0.077±0.003 | 0.473±0.003 | 3.260±0.010 |
| [16, 8, 4, 2, 2, 2] | [18, 10, 6, 4, 2, 2] | 0.058±0.003 | 0.490±0.002 | 3.168±0.009 |
| [16, 12, 4, 2, 2, 2] | [18, 16, 6, 4, 2, 2] | 0.044±0.004 | 0.508±0.003 | 3.063±0.010 |
| [16, 12, 6, 2, 2, 2] | [18, 16, 8, 4, 2, 2] | 0.038±0.002 | 0.515±0.003 | 2.951±0.008 |
We study the impact of Transformer depth and attention head configurations on HumanML3D (Table IV). Starting from a uniform setup ([4, 4, 4, 4, 4, 4] layers, [6, 6, 6, 6, 6, 6] heads), the model yields suboptimal results (FID 0.101, MM-Dist 3.359). As we adopt deeper, hierarchically structured layers and progressively narrower heads, performance improves consistently. The best configuration ([18, 16, 8, 4, 2, 2] layers, [16, 12, 6, 2, 2, 2] heads) achieves the lowest FID (0.079), highest R@1 (0.505), and lowest MM-Dist (3.002).
These findings suggest that allocating more depth in early stages and gradually reducing head width enables better modeling of both global semantics and local temporal details. Compared to uniformly scaled variants, hierarchical designs yield stronger R@1 and MM-Dist, highlighting the importance of architectural balance.
For more ablation details, please check our Appendix.
V Conclusion
@@
We presented MOGO, a one-pass autoregressive framework for high-quality and real-time 3D human motion generation. By combining the MoSA-VQ module for scale-adaptive residual vector quantization with the RQHC-Transformer for hierarchical causal decoding, MOGO achieves efficient and temporally coherent motion synthesis. The integration of TCA mechanism further enhances semantic alignment and enables better generalization in zero- and few-shot scenarios. Our framework supports low-latency, frame-by-frame inference, making it well-suited for real-time and interactive applications. Extensive experiments demonstrate that MOGO consistently outperforms prior transformer-based methods across multiple benchmarks in terms of generation quality and text-motion alignment, establishing a strong foundation for future research on controllable, editable, and interactive motion generation.
Appendix A Qualitative Results
A-A Qualitative Results of Infinite Motion Generation.
Figure 3 showcases qualitative results of our infinite motion generation pipeline. Each vertical dashed box represents a complete generation sequence conditioned on a given text prompt. From top to bottom, the first row shows the initial motion segment, while the subsequent rows illustrate recursively generated motion continuations, conditioned on the same or semantically extended instruction.
Across all four examples, our model demonstrates strong temporal consistency, with smooth transitions between motion segments and no noticeable discontinuities. The generated motions exhibit high semantic fidelity to the input descriptions, such as complex behaviors like “walks in a circle and then stop”, “jumps back and forth horizontally”, or “dancing forward with beats”, all of which are well-executed over extended sequences. The smooth transition between action phrases (e.g., “stands up and then walk in a circle”) further indicates that our model captures both short-term dynamics and long-horizon dependencies effectively.
To support reproducibility and further inspection, we provide all visual examples shown in Figure 3 as rendered videos, along with additional qualitative results in the supplementary materials. Specifically, the included .zip file contains not only video renderings corresponding to the figure, but also additional extended generations for diverse prompts—several of which feature single uninterrupted motion sequences exceeding 2000 frames in length. These results further demonstrate the stability of our model in synthesizing temporally consistent and semantically aligned motion over prolonged durations.
These results highlight the model’s capability to generate coherent, instruction-aligned motion indefinitely, making it suitable for real-world applications like virtual agents, gaming, and embodied language understanding.
A-B Qualitative Results of Single-Prompt Generation
Figure 4 presents a comprehensive overview of our model’s ability to generate motion sequences that align with a wide range of text prompts, demonstrating both semantic fidelity and fine-grained motion expressiveness. Each sample in the figure consists of a 3D motion mesh superimposed with motion trails (or motion blending) and its corresponding textual description, rendered in a stylized format where key action phrases are emphasized in red italics.
Semantic Diversity and Text-Condition Alignment.
Across the grid layout, we showcase 16 distinct prompts covering a diverse set of actions, such as:
Simple physical reactions: “The person was pushed but did not fall.”
Complex multi-phase motions: “A man is walking forward then steps over a stair then continues walking forward.”
Locomotion with intention shifts: “walked forward spun right on one foot and walked back to his original position.”
Fine-grained bodily behaviors: “holds their right foot with their right hand lifting both their left foot and left hand up.”
Stylized or culturally grounded motions: “The person does a salsa dance.”
Emotionally expressive behaviors: “A person walks with a limp, their right leg gets injured.”
These results indicate that our model is capable of capturing not only locomotion and spatial translation but also static poses, micro-movements, and non-trivial action transitions. The linguistic variations in the prompts are successfully translated into physically coherent movement, suggesting a high level of text-to-motion grounding.
Motion Coherence and Realism.
Visually, the rendered meshes show strong temporal consistency. The use of motion trails helps reveal the flow of the action over time, allowing readers to observe details such as:
Symmetrical limb movement (as in treadmill running),
Localized gestures (like jabbing with the hand),
Balanced posture shifts (e.g., squatting while crossing the leg and lowering with both hands).
Each figure shows a stable, physically plausible movement trajectory. Particularly in examples like the limp walking or the foot-holding pose, the model demonstrates a robust understanding of joint constraints and biomechanical plausibility, which are often hard to synthesize accurately.
Moreover, the transitions between joint configurations remain smooth and contextually aligned. For instance, the action “walks forward then steps over a stair” is depicted not only as a forward movement but with a discernible vertical leg lift, capturing the semantic intent of stepping.
Expressiveness and Subtle Behaviors.
The model also demonstrates the ability to synthesize expressive and subtle behaviors from relatively short prompts. Examples such as:
“blocks their face with their right arm”
“stands, crosses their left leg in front…”
highlight the model’s capability to generate contextually appropriate and biomechanically accurate poses. These motions require localized articulation and fine control, which the model executes effectively.
A-C Summary
In conclusion, Figure 4 provides compelling evidence of our model’s ability to generate temporally coherent, semantically grounded, and physically plausible motions across a wide range of instructions. The diversity of action types, from stylized dance to nuanced micro-movements, highlights its applicability to real-world embodied tasks such as virtual agents, assistive robots, and language-driven animation.
Appendix B Text Condition Alignment: Prompt-Engineering Strategy and Qualitative Study
B-A Prompt-Engineering Strategy
TCA relies on a single LLM call whose prompt is carefully engineered to yield: (i) a normalized, imperative restatement of the user request, (ii) a line-separated sequence of atomic motion steps, and (iii) a final rewritten instruction that merges all atomic actions into a fluent, single-sentence command with an explicit subject (e.g., “A person lifts both arms, swings them forward in an arc, and returns to a standing position”). The design combines several complementary techniques:
S1. Role & Goal Priming.
We begin with an explicit role declaration: “You are an instruction rewriter for a motion-generation robot.” This is immediately followed by a goal statement: “Transform the user request into a concise command (Normalized), a numbered list of executable steps (Steps), and a final rewritten instruction combining all actions into a single fluent sentence with a human subject.” This dual priming has two effects: (1) it reduces style drift across diverse user inputs, and (2) it biases the LLM toward a structured, imperative language style aligned with the motion decoder’s text prior.
S2. Chain-of-Thought (CoT) Cueing — Two-Stage Least-to-Most Prompting.
To obtain reliable, atomic motion commands without revealing the LLM’s private reasoning, we implement a lightweight Least-to-Most (LtM) procedure. The complete CoT [33] prompt is executed within a single API call but is conceptually divided into two internal stages:
S2.1 Intent Decomposition. The model is first instructed: “Think step by step. Rewrite the user request as one concise imperative sentence, then list the minimal executable actions, one per line, and finally output a single-sentence rewritten instruction beginning with ’A person …’ that combines all the actions. Do not reveal your reasoning.” This stage encourages hidden chain-of-thought while enforcing a deterministic, three-part output schema [34].
S2.2 Self-Verification & Refinement. Immediately after producing the draft list, the prompt appends: “Verify that every step can be executed in isolation by a humanoid motion agent. If any step is composite or ambiguous, split it further and rewrite the entire list and the final sentence accordingly.” The model then overwrites its own answer before the final stop token, yielding a refined set of atomic steps and a natural-language instruction that summarizes them with an explicit subject.
S3. Few-Shot Curriculum (In-Context Learning).
To equip the LLM with a prior on how free-form descriptions in our dataset should be reformulated into structured commands, we perform a lightweight in-context learning (ICL) setup. Specifically, we prepend three demonstration pairs to every query, sampled directly from our training corpus and rewritten into the target schema (Normalized + Steps + Final instruction). These demonstrations jointly span diverse [35] syntax—casual narration, imperative clause, and run-on sentence—and heterogeneous semantics—dance, locomotion, and object interaction. The curriculum is designed to:
expose the model to prompts consistent with the training data distribution, encouraging flexible segmentation into 2–5 atomic steps;
illustrate how varied tense and verbosity in raw captions can be normalized into concise imperative commands and rewritten human-centric instructions;
include both prompts with and without implicit pauses, teaching the model to insert explicit hold actions when necessary.
Action verbs in the demonstrations are drawn from the motion model’s training vocabulary, ensuring lexical consistency between the rewritten instructions and the downstream decoder, while requiring no parameter modifications. This ICL design narrows the style gap between descriptive captions in the dataset and real user prompts, improving controllability during TCA inference.
B-B Qualitative Examples
Example 1. (Humanml3D Test)
Raw prompt: “Someone imitating a golf swing.”
Normalized: “Raise arms, swing forward, return to stance.”
Thinking process:
1. Lift both arms as if holding a golf club
2. Swing the arms forward in an arc
3. Return to a relaxed standing posture -
Final rewritten instruction:
“A person lifts both arms, swings them forward in an arc, and returns to a standing position.”
Example 2. (Wild data Test)
Raw prompt: “A human is practicing expressive spinning gestures.”
Normalized: “Spin, pause, raise arms.”
Steps:
1. Spin once on the spot
2. Hold still for a beat
3. Raise both hands -
Final rewritten instruction:
“A person spins once on the spot, holds still briefly, and raises both hands.”
Example 3. (Wild data Test)
Raw prompt: “A human is rehearsing a dynamic forward motion sequence.”
Normalized: “Walk, crouch, roll forward, stand.”
Steps:
1. Walk forward
2. Drop into a crouch
3. Roll forward
4. Stand upright -
Final rewritten instruction:
“A person walks forward, drops into a crouch, rolls ahead, and stands upright.”
As shown in Figure 5, these cases show how TCA strategy reliably converts noisy, multi-clause requests into concise, executable motion scripts—enabling the decoder to render complex behaviours with accurate temporal structure without any motion-model fine-tuning.
Appendix C Ablation Study
C-A Impact of Learnable Scaling in MoSA-VQ.
@@ Appendix C Ablation Study
We investigate the role of the proposed learnable scaling mechanism in MoSA-VQ by comparing models with and without this design. In the baseline configuration, residual features at each quantization level are directly quantized without any adaptive normalization. In contrast, our full model applies a layer-specific affine transformation before quantization, allowing each level to adjust the residual space dynamically. As shown in Table V, incorporating learnable scale and bias parameters leads to consistent improvements across multiple evaluation metrics. The model demonstrates better reconstruction quality, enhanced alignment between motion and text, and stronger semantic consistency. These results highlight the benefit of introducing adaptive normalization in residual quantization, which facilitates more expressive and stable token representations.
w/ Learnable Scaling
| Codebook | FID | Top1 | MM-Dist |
@@
w/o Learnable Scaling
| Codebook | FID | Top1 | MM-Dist |
@@
C-B Impact of prompt and quantization.
C-B1 Without PnQ
An additional layer token is inserted between the textual prompt and the quantized motion sequence, serving as an intermediate fusion node. However, this design introduces a level of indirection that leads to redundant attention computation and diluted gradient flow. As a result, the model exhibits weaker cross-modal alignment, as the semantic linkage between text and motion must be learned indirectly through this intermediate token, which may act as a bottleneck in representation transfer.
C-B2 With PnQ
We instead fuse the prompt token directly with the quantization condition tokens through positional and conditional injection. This eliminates the need for an intermediate token and allows the model to establish a more direct and semantically grounded connection between the text prompt and motion representation. Consequently, this design not only improves attention efficiency—by reducing unnecessary detours in the attention graph—but also enhances the cross-modal alignment by preserving semantic consistency throughout the sequence.
C-C Effect of Dataset Size
(a) Dataset Size vs. FID
| Epochs | 0% CMP | 50% CMP | 100% CMP |
| 50 | 0.671 | 0.601 | 0.501 |
| 80 | 0.437 | 0.379 | 0.284 |
| 100 | 0.347 | 0.263 | 0.205 |
(b) Real-time Inference (A800)
| Method | AITS | Params | FID |
| MLD | 0.14 | 34M | 0.473 |
| MotionGPT | 0.81 | 770M | 0.160 |
| MoMask | 0.06 | 163M | 0.045 |
| MOGO | 0.37 | 830M | 0.038 |
To investigate the scalability of MOGO, we evaluate its performance under different training set sizes by progressively mixing HumanML3D with 50% and 100% proportions of the CMP dataset. Crucially, we keep the MoSA-VQ quantizer frozen during this process to isolate the impact of data volume on the downstream generation quality.
As shown in Table VI(a), increasing the amount of training data leads to consistent improvements in FID across training epochs. For instance, at 100 training epochs, FID drops from 0.347 (with only HumanML3D) to 0.263 with 50% CMP, and further to 0.205 with full CMP integration. This trend holds across earlier epochs as well, demonstrating that larger and more diverse motion data helps the model generalize better, even without modifying the discrete representation space.
These results highlight the scalability and data efficiency of MOGO’s architecture, as well as the robustness of the MoSA-VQ quantizer in accommodating novel distributions through fixed codes.
C-D Real-Time Inference Efficiency
We evaluate the inference efficiency of MOGO by measuring the average time required to generate a full motion sequence of 196 frames per input sentence, denoted as Average Inference Time per Sentence (AITS). As shown in Table VI(b), we compare our model with several state-of-the-art methods under consistent hardware (single A800 GPU).
MOGO achieves an average inference time of 0.37 seconds per sentence, demonstrating its ability to generate long-form motion in near real-time. While MoMask is faster due to its significantly smaller architecture, it delivers inferior performance in motion quality compared to MOGO. Conversely, methods like MotionGPT and MLD either trade off fidelity or speed, with MLD exceeding typical latency thresholds for interactive applications.
Despite its large parameter size (830M), MOGO produces the highest motion quality (FID = 0.038) among all methods while maintaining a latency level that is well-suited for real-time deployment in applications such as virtual avatars, embodied agents, and interactive character control.
These results confirm that MOGO is not only state-of-the-art in generation quality, but also practically deployable in real-time or near-real-time environments.
C-E Effect of Codebook Size